A container ship with a displacement of 15,200 T and is sailing at 20 knots and comes to a stop after 0.75h. What is the work done on the container ship if it covers a stopping distance of 3,125 m?


1
Expert's answer
2021-11-01T12:54:44-0400

Solution.

"m=15 200 T=15200000kg;"

"t=0.75h=2700s;"

"v_0=20knots=10.29 m\/s;"

"v=0;"

"l=3125m;"

"W=Fl;"

"F=ma;"

"a=\dfrac{v-v_0}{t};" "a=\dfrac{0-10.29}{2700}=-0.0038 m/s^2;"

"F=15200000\sdot ( -0.0038)=-57928.9N;"

"W=-57928.9\sdot 3125=-181027778J=-181MJ;"

Answer: "W=-181MJ."